### v3

@@ 4.1 RMU on h-subproblem and W-subproblem
Figure 2: Left: A typical convergence of the four methods: EPG, RMU, RADMM and RALM. In the y-axis, $F_{\min}$ refers to the smallest achieved $F$ value across the methods. The smaller subplot is zooming the first $7\times 10^{-3}$ seconds. In this test, we have $(m,r)=(5,3)$ and RMU is the fastest methods with strict feasibility. Right: the results on different settings of $(m,r)$. A ranking vector $(x,y,z,...)$ means the method has $x$ times out of 100 runs having the best performance, $y$ times out of 100 runs having the 2nd performance, and so on. There are four methods, and the fifth number indicate how many times the method produces an infeasible solution ( having negative values) in the end.
@@ 4.2 Testing Chordal-NMF on synthetic dataset
Synthetic Dataset. We use a specific dataset $\bm{M}^{\epsilon,\delta}=\bm{W}_{\text{true}}\bm{H}_{\text{true}}^{\epsilon,\delta}$ with $\epsilon>0,\delta>0$ as
@@ 4.3 Testing Chordal-NMF on real-world data
Further analysis on the three selected areas. In Figure. 5, three boxes were chosen to represent three distinct cloud conditions: the red box is a cloudy area, the cyan box is a less-cloudy areas, and the yellow box is a cloudless area. We compute the spectral signatures of the pixels in these areas, and their mean behavior is plotted in Figs. 6. We numerically compared these spectral profile vectors by two criteria:
@@
[…]
    $\text{SID}=\sum\limits_{i=1}^{N}{p_{i}\log{\frac{p_{i}}{q_{i}}}}+\sum\limits_{i=1}^{N}{q_{i}\log{\frac{q_{i}}{p_{i}}}},\quad\alpha=\cos^{-1}\left(\sum\limits_{i=1}^{N}{t_{i}r_{i}}\Big{/}\sqrt{\bigg{(}\sum\limits_{i=1}^{N}{t_{i}^{2}}\bigg{)}\bigg{(}\sum\limits_{i=1}^{N}{r_{i}^{2}}\bigg{)}}\right).$
@@ Appendix A The proof of Proposition 1
$\begin{array}[]{rcl}\nabla f(\bm{x})&=&\displaystyle\dfrac{\|\bm{D}\bm{x}+\bm{e}\|_{2}}{\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle}\dfrac{\|\bm{D}\bm{x}+\bm{e}\|_{2}^{2}\nabla\Big{(}\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\Big{)}^{2}-\Big{(}\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\Big{)}^{2}\nabla\|\bm{D}\bm{x}+\bm{e}\|_{2}^{2}}{2\|\bm{D}\bm{x}+\bm{e}\|_{2}^{4}}\vspace{0.5mm}\\ &=&\displaystyle\dfrac{\|\bm{D}\bm{x}+\bm{e}\|_{2}^{2}\nabla\Big{(}\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\Big{)}^{2}-\Big{(}\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\Big{)}^{2}\nabla\|\bm{D}\bm{x}+\bm{e}\|_{2}^{2}}{2\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\|\bm{D}\bm{x}+\bm{e}\|_{2}^{3}}\vspace{0.5mm}\\ &=&\displaystyle\dfrac{\|\bm{D}\bm{x}+\bm{e}\|_{2}^{2}2\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\bm{A}^{\top}\bm{c}-\Big{(}\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\Big{)}^{2}2\bm{D}^{\top}(\bm{D}\bm{x}+\bm{e})}{2\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\|\bm{D}\bm{x}+\bm{e}\|_{2}^{3}}\vspace{0.5mm}\\ &=&\displaystyle\dfrac{\|\bm{D}\bm{x}+\bm{e}\|_{2}^{2}\bm{A}^{\top}\bm{c}-\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\bm{D}^{\top}(\bm{D}\bm{x}+\bm{e})}{\|\bm{D}\bm{x}+\bm{e}\|_{2}^{3}}.\end{array}$
@@ Appendix C Derivation of content in Table 1
Consider (W-subproblem) with $n=1$. Following the discussion in §3.1, we get grid of the denominator $\langle\bm{W},\bm{W}\bm{A}_{1}\rangle_{\textrm{F}}$ in by introducing a constrained problem $\operatorname*{\textrm{argmax}\,}_{\bm{W}\geq\bm{0}}\langle\bm{B}_{j},\bm{W}\rangle_{\textrm{F}}\,\,\textrm{s.t.}\,\,\langle\bm{W},\bm{W}\bm{A}_{1}\rangle_{\textrm{F}}=1$. We note that $\bm{A}_{1}$ is a rank-1 symmetric positive semi-definite (PSD) matrix with two eigenvalues: a single positive eigenvalue and 00 with multiplicity $r-1$. Moreover, $\bm{A}_{1}$ has its square-root $\bm{A}^{1/2}_{1}$, so $\langle\bm{W}^{\top}\bm{W},\bm{A}_{1}\rangle_{\textrm{F}}=\langle\bm{W}^{\top}\bm{W},\bm{A}^{1/2}_{1}\bm{A}^{1/2}_{1}\rangle_{\textrm{F}}=\langle\bm{W}\bm{A}^{1/2}_{1},\bm{W}\bm{A}^{1/2}_{1}\rangle_{\textrm{F}}=\langle\bm{W},\bm{W}\rangle_{\bm{A}^{1/2}_{1}}$ allows us to define the manifold $\mathcal{M}_{1}$ as
@@
${\mathcal{R}_{\bm{Z}}(\bm{W})=(\bm{W}+\bm{Z})/\|\bm{Z}+\bm{W}\|_{\bm{A}_{j}^{1/2}}\overset{T_{\bm{Z}}\mathcal{M}_{j}}{=}(\bm{Z}+\bm{W})/\sqrt{1+\|\bm{W}\|_{\bm{A}_{j}^{1/2}}^{2}}},$ ($\mathcal{R}$)
@@ Appendix D Details of manifold approach for W-subproblem with $n=2$
$\textrm{D}h(\bm{W})^{*}[{\bm{\alpha}}]=2\alpha_{1}\bm{W}\bm{A}_{1}+2\alpha_{2}\bm{W}\bm{A}_{2}=2\bm{W}(\alpha_{1}\bm{A}_{1}+\alpha_{2}\bm{A}_{2}).$
@@
$\bm{S}_{\bm{Z}}^{\top}\bm{S}_{\bm{Z}}\begin{bmatrix}\alpha_{1}^{*}\\ \alpha_{2}^{*}\end{bmatrix}~{}=~{}\dfrac{1}{2}\bm{S}_{\bm{Z}}^{\top}\textrm{vec}\bm{W}~{}~{}\implies~{}~{}\bm{\alpha}^{*}=\dfrac{1}{2}(\bm{S}_{\bm{Z}}^{\top}\bm{S}_{\bm{Z}})^{-1}\bm{S}_{\bm{Z}}^{\top}\textrm{vec}\bm{W}=\dfrac{1}{2}\bm{S}_{\bm{Z}}^{\dagger}\textrm{vec}\bm{W}.$
@@
$\textrm{proj}_{T_{\bm{Z}}\mathcal{M}^{[2]}}(\bm{W})=\bm{W}-2\bm{Z}\sum_{j=1}^{2}\dfrac{1}{2}\Big{(}\bm{S}_{\bm{Z}}^{\dagger}\textrm{vec}\bm{W}\Big{)}_{j}\bm{A}_{j}=\bm{W}-2\bm{Z}(\alpha_{1}^{*}\bm{A}_{1}+\alpha_{2}^{*}\bm{A}_{2}).$ (7)
@@
Note that the value of $\alpha_{1}^{*}$ and $\alpha_{2}^{*}$ in (8) is an implicit function of $\bm{Z}$ and $\nabla F(\bm{W})$ as we have that $\begin{bmatrix}\alpha_{1}^{*}\\ \alpha_{2}^{*}\end{bmatrix}=\dfrac{1}{2}\begin{bmatrix}\Big{(}\bm{S}_{\bm{Z}}^{\dagger}\textrm{vec}\nabla F(\bm{W})\Big{)}_{1}\\ \Big{(}\bm{S}_{\bm{Z}}^{\dagger}\textrm{vec}\nabla F(\bm{W})\Big{)}_{2}\end{bmatrix}$ So the explicit expression of the Riemannian gradient is
$\textrm{grad}F(\bm{W})=\nabla F(\bm{W})-\bm{Z}\Big{(}\bm{S}_{\bm{Z}}^{\dagger}\textrm{vec}\nabla F(\bm{W})\Big{)}_{1}\bm{A}_{1}-\bm{Z}\Big{(}\bm{S}_{\bm{Z}}^{\dagger}\textrm{vec}\nabla F(\bm{W})\Big{)}_{2}\bm{A}_{2}.$

### v4

@@ 4.1 RMU on h-subproblem and W-subproblem
Figure 2: Left: A typical convergence of the four methods: EPG, RMU, RADMM and RALM. In the y-axis, $F_{\min}$ refers to the smallest achieved $F$ value across the methods. The smaller subplot is zooming the first $7\times 10^{-3}$ seconds. In this test, we have $(m,r)=(5,3)$ and RMU is the fastest methods with strict feasibility. Right: the results on different settings of $(m,r)$. A ranking vector (x,y,z,)(x,y,z,...) means the method has $x$ times out of 100 runs having the best performance, $y$ times out of 100 runs having the 2nd performance, and so on. There are four methods, and the fifth number indicate how many times the method produces an infeasible solution ( having negative values) in the end.
@@ 4.2 Testing Chordal-NMF on synthetic dataset
Synthetic dataset. We use a dataset $\bm{M}^{\epsilon,\delta}=\bm{W}_{\text{true}}\bm{H}_{\text{true}}^{\epsilon,\delta}$ with $\epsilon>0,\delta>0$ as
@@ 4.3 Testing Chordal-NMF on real-world data
Further analysis on the three selected areas. In Fig. 5, three boxes were chosen to represent three distinct cloud conditions: the red box is a cloudy area, the cyan box is a less-cloudy areas, and the yellow box is a cloudless area. We compute the spectral signatures of the pixels in these areas, and their mean behavior is plotted in Fig. 6. We numerically compared these spectral profile vectors by two criteria:
@@
    SID-SAM between the spectral profile vectors. Given two spectral profile vectors $\bm{t}=[t_{1},...,t_{N}]$ and $\bm{r}=[r_{1},...,r_{N}]$, we compute $\bm{p}=\bm{t}/\|\bm{t}\|_{2}$, $\bm{q}=\bm{r}/\|\bm{r}\|_{2}$, and then we compute the SID-SAM [28] of the spectral information divergence (SID) and the spectral angle mapper (SAM), defined as $\text{SID-SAM}=\text{SID}\times\tan{\alpha}$, where
    $\text{SID}=\sum\limits_{i=1}^{N}{p_{i}\log{\frac{p_{i}}{q_{i}}}}+\sum\limits_{i=1}^{N}{q_{i}\log{\frac{q_{i}}{p_{i}}}},\quad\alpha=\cos^{-1}\left(\sum\limits_{i=1}^{N}{t_{i}r_{i}}\bigg/\sqrt{\bigg(\sum\limits_{i=1}^{N}{t_{i}^{2}}\bigg)\bigg(\sum\limits_{i=1}^{N}{r_{i}^{2}}\bigg)}\right).$
@@ Appendix A The proof of Proposition 1
$\begin{array}[]{rcl}\nabla f(\bm{x})&=&\displaystyle\dfrac{\|\bm{D}\bm{x}+\bm{e}\|_{2}}{\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle}\dfrac{\|\bm{D}\bm{x}+\bm{e}\|_{2}^{2}\nabla\big(\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\big)^{2}-\big(\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\big)^{2}\nabla\|\bm{D}\bm{x}+\bm{e}\|_{2}^{2}}{2\|\bm{D}\bm{x}+\bm{e}\|_{2}^{4}}\vskip 1.42262pt\\ &=&\displaystyle\dfrac{\|\bm{D}\bm{x}+\bm{e}\|_{2}^{2}\nabla\big(\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\big)^{2}-\big(\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\big)^{2}\nabla\|\bm{D}\bm{x}+\bm{e}\|_{2}^{2}}{2\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\|\bm{D}\bm{x}+\bm{e}\|_{2}^{3}}\vskip 1.42262pt\\ &=&\displaystyle\dfrac{\|\bm{D}\bm{x}+\bm{e}\|_{2}^{2}2\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\bm{A}^{\top}\bm{c}-\big(\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\big)^{2}2\bm{D}^{\top}(\bm{D}\bm{x}+\bm{e})}{2\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\|\bm{D}\bm{x}+\bm{e}\|_{2}^{3}}\vskip 1.42262pt\\ &=&\displaystyle\dfrac{\|\bm{D}\bm{x}+\bm{e}\|_{2}^{2}\bm{A}^{\top}\bm{c}-\langle\bm{A}\bm{x}+\bm{b},\bm{c}\rangle\bm{D}^{\top}(\bm{D}\bm{x}+\bm{e})}{\|\bm{D}\bm{x}+\bm{e}\|_{2}^{3}}.\end{array}$
@@ Appendix C Derivation of content in Table 1
Consider (W-subproblem) with $n=1$. Following the discussion in §3.1, we get grid of the denominator $\langle\bm{W},\bm{W}\bm{A}_{1}\rangle_{\textrm{F}}$ in by introducing a constrained problem $\operatorname*{\textrm{argmax}\,}_{\bm{W}\geq\bm{0}}\langle\bm{B}_{j},\bm{W}\rangle_{\textrm{F}}\,\,\textrm{s.t.}\,\,\langle\bm{W},\bm{W}\bm{A}_{1}\rangle_{\textrm{F}}=1$. We note that $\bm{A}_{1}$ is a rank-1 symmetric positive semi-definite (PSD) matrix with two eigenvalues: a single positive eigenvalue and 0 with multiplicity r1r-1. Moreover, $\bm{A}_{1}$ has its square-root $\bm{A}^{1/2}_{1}$, so $\langle\bm{W}^{\top}\bm{W},\bm{A}_{1}\rangle_{\textrm{F}}=\langle\bm{W}^{\top}\bm{W},\bm{A}^{1/2}_{1}\bm{A}^{1/2}_{1}\rangle_{\textrm{F}}=\langle\bm{W}\bm{A}^{1/2}_{1},\bm{W}\bm{A}^{1/2}_{1}\rangle_{\textrm{F}}=\langle\bm{W},\bm{W}\rangle_{\bm{A}^{1/2}_{1}}$ allows us to define the manifold $\mathcal{M}_{1}$ as
@@
$\mathcal{R}_{\bm{Z}}(\bm{W})~=~\dfrac{\bm{W}+\bm{Z}}{\|\bm{Z}+\bm{W}\|_{\bm{A}_{j}^{1/2}}}\overset{T_{\bm{Z}}\mathcal{M}_{j}}{=}\dfrac{\bm{Z}+\bm{W}}{\sqrt{1+\|\bm{W}\|_{\bm{A}_{j}^{1/2}}^{2}}},$ ($\mathcal{R}$)
@@ Appendix D Details of manifold approach for W-subproblem with $n=2$
$\textrm{D}h(\bm{W})^{*}[{\bm{\alpha}}]~=~2\alpha_{1}\bm{W}\bm{A}_{1}+2\alpha_{2}\bm{W}\bm{A}_{2}~=~2\bm{W}(\alpha_{1}\bm{A}_{1}+\alpha_{2}\bm{A}_{2}).$
@@
$\bm{S}_{\bm{Z}}^{\top}\bm{S}_{\bm{Z}}\begin{bmatrix}\alpha_{1}^{*}\\ \alpha_{2}^{*}\end{bmatrix}~=~\dfrac{1}{2}\bm{S}_{\bm{Z}}^{\top}\textrm{vec}\bm{W}~~\implies~~\bm{\alpha}^{*}~=~\dfrac{1}{2}(\bm{S}_{\bm{Z}}^{\top}\bm{S}_{\bm{Z}})^{-1}\bm{S}_{\bm{Z}}^{\top}\textrm{vec}\bm{W}~=~\dfrac{1}{2}\bm{S}_{\bm{Z}}^{\dagger}\textrm{vec}\bm{W}.$
@@
$\textrm{proj}_{T_{\bm{Z}}\mathcal{M}^{[2]}}(\bm{W})~=~\bm{W}-2\bm{Z}\sum_{j=1}^{2}\dfrac{1}{2}\Big(\bm{S}_{\bm{Z}}^{\dagger}\textrm{vec}\bm{W}\Big)_{j}\bm{A}_{j}~=~\bm{W}-2\bm{Z}(\alpha_{1}^{*}\bm{A}_{1}+\alpha_{2}^{*}\bm{A}_{2}).$ (7)
@@
Note that the value of $\alpha_{1}^{*}$ and $\alpha_{2}^{*}$ in (8) is an implicit function of $\bm{Z}$ and $\nabla F(\bm{W})$ as we have that $\begin{bmatrix}\alpha_{1}^{*}\\ \alpha_{2}^{*}\end{bmatrix}=\dfrac{1}{2}\begin{bmatrix}\Big(\bm{S}_{\bm{Z}}^{\dagger}\textrm{vec}\nabla F(\bm{W})\Big)_{1}\\ \Big(\bm{S}_{\bm{Z}}^{\dagger}\textrm{vec}\nabla F(\bm{W})\Big)_{2}\end{bmatrix}$ . So the explicit expression of the Riemannian gradient is
$\textrm{grad}F(\bm{W})~=~\nabla F(\bm{W})-\bm{Z}\Big(\bm{S}_{\bm{Z}}^{\dagger}\textrm{vec}\nabla F(\bm{W})\Big)_{1}\bm{A}_{1}-\bm{Z}\Big(\bm{S}_{\bm{Z}}^{\dagger}\textrm{vec}\nabla F(\bm{W})\Big)_{2}\bm{A}_{2}.$